OpenStudy (anonymous):

The number of cars at a dealership dropped from 64 to 48 after a weekend sale. What is the percent of decrease in the number of cars? A. 20% B. 25% C. 33% D. 75%

OpenStudy (gorv):

decrease=64-48=16 %decreaswe=(16/64)*100=(1/4)*100=25%

OpenStudy (texaschic101):

percent decrease = (older - newer) / older percent decrease = (64 - 48) / 64 percent decrease = 16/64 = .25 turned to a percent = 25%

OpenStudy (anonymous):

The price of a bracelet increased from $170 to $200 on Mother's Day. What is the percent of increase in the price? Round to the nearest percent. A. 15% B. 18% C. 30% D. 85%

OpenStudy (anonymous):

I got B, what usually do is just multiply the %'s by the number in this case 170

OpenStudy (anonymous):

The number of dogs in a pet store decreased from 35 to 31. What is the percent of decrease in the number of dogs? Round to the nearest percent. A. 11% B. 13% C. 40% D. 97%

OpenStudy (anonymous):

Dude you now know how to work these out now... We're are not going to d your homework

OpenStudy (texaschic101):

percent increase = (new - old) / old percent increase = (200 - 170) / 170 percent increase = 30/170 = 0.176 rounds to 0.18, turned to percent = 18%

OpenStudy (texaschic101):

now set the last problem up just like the first problem. You can do this.

OpenStudy (anonymous):

A or B

OpenStudy (anonymous):

?